Grade 11General Physics

the height from the surface of earth at which the total energy of the satellite is equal to its potential energy at a height 2R from the surface of earth is (R= radius of the earth)

ranganathan
8 Years agoGrade 11

1 Answer

Nani yeggireddy
8 Years ago
Total energy = - GMm/2(R+h)Potential energy at height 2R = - GMm/(R+2R) = - GMm/3RGiven : Total energy = potential energy at height 2R -GMm/2(R+h) = - GMm/3R When, h=R/2-GMm/2(R+R/2)= - GMm/3R -GMm/2(3R/2) = - GMm /3R -GMm /3R = -GMm /3R Therefore at height R/2 the total energy of the satellite is equal to potential energy at height 2R..